# Learning Theory for Kernel Bilevel Optimization

### v1

@@ Appendix E Maximal Inequalities for Bounded and Lipschitz Family of Functions
For a given probability measure $\mu$, denote by $L_{2}(\mu)$ the space of square $\mu$-integrable real-valued functions. We denote by ${\left\lVert f\right\rVert}_{\mathbb{D},2}\coloneqq\mathbb{E}_{\mathbb{D}}{\left[f(z)^{2}\right]}^{\frac{1}{2}}$ the $L_{2}(\mu)$-norm of any function $f\in L_{2}(\mu)$. For any $\epsilon>0$, we denote by $D\left(\epsilon,\mathcal{T},L_{2}(\mu)\right)$ the $\epsilon$-packing number of $\mathcal{T}$ w.r.t. $L_{2}(\mu)$. The next proposition provides a control on such a number under regularity conditions on the family $\mathcal{T}$.
Proposition E.1 (Control on the packing number).
Assume that $\Omega$ is a compact subset of $\mathbb{R}^{d}$, that the parametric family $\mathcal{T}$ defined in Equation 19 is uniformly bounded by a positive constant $M$, and that there exists a positive constant $L$ so that, for any measure $\mu$, $\omega\mapsto t_{\omega}(z)$ is $L$-Lipschitz for any $z\in\mathcal{Z}$. Then, there exists a positive constant $c(\Omega)$ greater than $1$ that depends only on $\Omega$ and $d$ so that, for any probability measure $\mu$ on $\mathcal{Z}$, the following bound holds for any $0<\epsilon\leq M$:
@@
First using (Kosorok, 2008, Lemma 9.18) and (Kosorok, 2008, Paragraph 8.1.2), we know that the $\epsilon$-packing number $D\left(\epsilon,\mathcal{T},L_{2}(\mu)\right)$ is smaller than the $\frac{\epsilon}{2}$-bracketing number $N_{[]}\left(\frac{\epsilon}{2},\mathcal{T},L_{2}(\mu)\right)$. Hence, we only need to control the bracketing number. To this end, we recall that the function $\omega\mapsto t_{\omega}(z)$ is $L$-Lipschitz for any $z\in\mathcal{Z}$, so that (Van der Vaart, 2000, Example 19.7) ensures the existence of a positive constant $c(\Omega)$ that depends only on $\Omega$ for which the following inequality holds for any $0<\epsilon<L\text{diam}(\Omega)$:
@@
Theorem E.2 (Maximal inequality for degenerate bounded and Lipschitz $U$-processes).
Let $k$ be either $1$ or $2$. Consider a parametric family $\mathcal{T}\coloneqq{\left\{\otimes^{k}\mathcal{Z}\ni(z_{1},\ldots,z_{k})\mapsto t_{\omega}(z_{1},\ldots,z_{k})\in\mathbb{R}\mid\omega\in\Omega\right\}}$ of real-valued functions over $\otimes^{k}\mathcal{Z}$ and indexed by a parameter $\omega\in\Omega$, where $\Omega$ is a compact subset of $\mathbb{R}^{d}$. For a given probability distribution $\mathbb{D}$ over $\mathcal{Z}$, assume that all elements $t_{\omega}$ are degenerate w.r.t. $\mathbb{D}$, meaning that:
@@
$U_{n}^{k}t_{\omega}\coloneqq\begin{cases}\frac{1}{n}\sum_{i=1}^{n}t_{\omega}(z_{i}),&\qquad\text{if}\quad k=1\\ \frac{1}{n(n-1)}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}t_{\omega}(z_{i},z_{j}),&\qquad\text{if}\quad k=2.\end{cases}$
@@
Maximal inequality for degenerate $U$-processes. We will first apply the general result in (Sherman, 1994, Maximal inequality) which controls $\mathbb{E}_{\mathbb{D}}{\left[\sup_{\omega\in\Omega}{\left\lvert U_{n}^{k}t_{\omega}\right\rvert}\right]}$ in terms of the packing number of $\mathcal{T}$. First note, by assumption, that the functions $t_{\omega}(z_{1},\ldots,z_{k})$ are uniformly bounded by a positive constant $M$. Therefore, the constant function $T(z_{1},\ldots,z_{k})\coloneqq M$ is an envelope for $\mathcal{T}$, i.e., $T$ satisfies $T(z_{1},\ldots,z_{k})\geq\sup_{\omega\in\Omega}{\left\lvert t_{\omega}(z_{1},\ldots,z_{k})\right\rvert}$ for any $(z_{1},\ldots,z_{k})\in\otimes^{k}\mathcal{Z}$. The envelope $T$ is, a fortiori, square $\mu$-integrable for any probability measure $\mu$ on $\otimes^{k}\mathcal{Z}$. Hence, we can apply (Sherman, 1994, Maximal inequality) with the choice $T$ for the envelope function and set the integer $m$ appearing in the result to $m=d$ to get the following bound:
@@
where $\Gamma$ is a positive universal constant [^2] The constant $\Gamma$ appearing (Sherman, 1994, Maximal inequality) depends only on $k$, $p$ and $m$, i.e., $\Gamma\coloneqq g(k,p,m)$. Since, we are only interested in $k\leq 2$ and $p\leq 2$ and $m$ is fixed to $d$, we choose $\Gamma$ to be $\max_{1\leq k,p\leq 2}g(k,p,d)^{\frac{1}{p}}$, so that it is the same in all our cases. that depends only on $d$ and that we choose to be greater than $1$, while $\mu_{n}$ are suitably chosen probability measures on $\otimes^{k}\mathcal{Z}$ that possibly depend on the samples $z_{1},\ldots,z_{n}$ and other random variables, and $\delta_{n}{\left\lVert T\right\rVert}_{\mu_{n},2}\coloneqq\sup_{\omega\in\Omega}{\left\lVert t_{\omega}\right\rVert}_{\mu_{n},2}$. Here, the expectation symbol in the right-hand side is over all randomness on which $\mu_{n}$ might depend. Note that the original result in (Sherman, 1994, Maximal inequality) is stated using a slightly different definition of the packing number but which is still equivalent to the statement above in our setting [^3] In (Sherman, 1994, Maximal inequality), the author considers a modified version of the $\epsilon$-packing number (call it $\tilde{D}(\epsilon,\mathcal{T},L_{2}(\mu))$) associated to $L_{2}(\mu)$ but endowed with a normalized version of the standard norm on $L_{2}(\mu)$: ${\left\lVert f\right\rVert}_{\mu}\coloneqq\frac{{\left\lVert f\right\rVert}_{\mu,2}}{{\left\lVert T\right\rVert}_{\mu,2}}$. Both numbers are related by the following identity: $\tilde{D}{\left(\epsilon,\mathcal{T},L_{2}(\mu)\right)}=D(\epsilon{{\left\lVert T\right\rVert}_{\mu,2}},\mathcal{T},L_{2}(\mu))$, thus making the statement (20) equivalent to the original statement in (Sherman, 1994, Maximal inequality). .
In our setting, the envelope function is constant and equal to $M$, and by definition $\delta_{n}\leq 1$. Hence, the inequality in Equation 20 further becomes:
$\mathbb{E}_{\mathbb{D}}{\left[\sup_{\omega\in\Omega}{\left\lvert U_{n}^{k}t_{\omega}\right\rvert}^{p}\right]}^{\frac{1}{p}}\leq n^{-\frac{k}{2}}M\Gamma\mathbb{E}_{\mathbb{D}}{\left[\int_{0}^{1}{\left(D{\left(\epsilon{\left\lVert T\right\rVert}_{\mu_{n},2},\mathcal{T},L_{2}(\mu_{n})\right)}\right)}^{\frac{1}{2dp}}\mathop{}\!\mathrm{d}\epsilon\right]}.$ (21)
@@
Control on the packing number. We have shown that the constant function $T(z_{1},\ldots,z_{k})\coloneqq M$ is an envelope for $\mathcal{T}$ which is, a fortiori, square $\mu$-integrable for any probability measure $\mu$ with ${\left\lVert T\right\rVert}_{\mu,2}=M<+\infty$. Moreover, the functions $\omega\mapsto t_{\omega}(z_{1},\ldots,z_{k})$ are $L$-Lipschitz for any $(z_{1},\ldots,z_{k})\in\otimes^{k}\mathcal{Z}$. We can therefore apply Proposition E.1 which ensures the existence of a positive constant $c(\Omega)$ greater than $1$ and that depends only on $\Omega$ and $d$ so that the following estimate on the $\epsilon$-packing number of the class $\mathcal{T}$ w.r.t. $L_{2}(\mu)$ holds:
@@
Combining Equation 22 with Equation 21 yields:
$\mathbb{E}_{\mathbb{D}}{\left[\sup_{\omega\in\Omega}{\left\lvert U_{n}^{k}t_{\omega}\right\rvert}^{p}\right]}^{\frac{1}{p}}$ $\leq n^{-\frac{k}{2}}M\Gamma\mathbb{E}_{\mathbb{D}}{\left[\int_{0}^{1}{\left(A\epsilon^{-d}\right)}^{\frac{1}{2dp}}\mathop{}\!\mathrm{d}\epsilon\right]}=n^{-\frac{k}{2}}M\Gamma A^{\frac{1}{2dp}}\underbrace{\int_{0}^{1}\epsilon^{-\frac{1}{2p}}\mathop{}\!\mathrm{d}\epsilon}_{\leq 2}$
@@
The following next propositions are particular instances of Theorem E.2 and will be used to obtain the main bounds.
Proposition E.3 (Maximal inequality for empirical processes).
@@
The upper-bound is a direct consequence of Theorem E.2. Indeed consider the family $\mathcal{S}$ of functions of the form $s_{\omega}(z)=t_{\omega}(z)-\mathbb{E}_{\bar{z}\sim\mathbb{D}}{\left[t_{\omega}(\bar{z})\right]}$, for any $z\in\mathcal{Z}$. Then clearly, the process $U_{n}^{1}s_{\omega}\coloneqq\frac{1}{n}\sum_{i=1}^{n}s_{\omega}(z_{i})$ is degenerate of order $k=1$, and the family $\mathcal{S}$ is uniformly bounded by $2M$ and is $2L$-Lipschitz. Hence, by Theorem E.2, the following maximal inequality holds:
@@
We get the desired upper-bound by redefining $c(\Omega)$ to contain the factor $2$. ∎
Proposition E.4 (Maximal inequality for $U$-processes of order 2).
@@
The proof will proceed by first decomposing $\tau_{\omega}$ into a sum of a degenerate $U$-process and a term of order $O(\frac{1}{n})$. The maximal inequality for degenerate $U$-processes from (Sherman, 1994) will be employed to obtain the desired bound.
@@
$s_{\omega}(z,z^{\prime})=t_{\omega}(z,z^{\prime})-\mathbb{E}_{\bar{z}\sim\mathbb{D}}\left[t_{\omega}(z,\bar{z})\right]-\mathbb{E}_{\bar{z}\sim\mathbb{D}}\left[t_{\omega}(\bar{z},z^{\prime})\right]+\mathbb{E}_{\{\bar{z},\underline{z}\}\sim\mathbb{D}\otimes\mathbb{D}}\left[t_{\omega}(\bar{z},\underline{z})\right].$ (23)
@@
$\tau_{\omega}=\overbrace{\frac{1}{(n-1)}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}s_{\omega}\big{(}z_{i},z_{j}\big{)}}^{U_{n}^{2}s_{\omega}}-\frac{1}{n^{2}(n-1)}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}t_{\omega}(z_{i},z_{j})+\frac{1}{n^{2}}\sum_{i=1}^{n}t_{\omega}(z_{i},z_{i}).$
@@
Hence, it only remains to control the first term in the above inequality. To this end, we will use a maximal inequality for degenerate $U$-processes due to (Sherman, 1994).
Maximal inequality for degenerate $U$-processes. We will first check that $U_{n}^{2}s_{\omega}$ is a degenerate statistic for a given $\omega\in\Omega$. To this end, simple calculations show that for any $z$ in $\mathcal{Z}$:
@@
The above equalities precisely ensure that $U_{n}^{2}s_{\omega}$ is a degenerate $U$-statistic for $\mathbb{D}$. Consider now the family $\mathcal{S}\coloneqq{\left\{\mathcal{Z}\times\mathcal{Z}\ni(z,z^{\prime})\mapsto s_{\omega}(z,z^{\prime})\in\mathbb{R}\mid\omega\in\Omega\right\}}$. We show that $\mathcal{S}$ is uniformly bounded and Lipschitz which allows to directly apply the result stated in Theorem E.2, which is a special case of the more general result in (Sherman, 1994, Maximal inequality). First note, by assumption, that the functions $t_{\omega}(z,z^{\prime})$ are uniformly bounded by a positive constant $M$. Hence, using Equation 23, it follows that $s_{\omega}(z,z^{\prime})$ is uniformly bounded by $4M$. Moreover, the functions $\omega\mapsto t_{\omega}(z,z^{\prime})$ are $L$-Lipschitz for any $z,z^{\prime}$ in $\mathcal{Z}$. Hence, from Equation 23, we directly have that $\omega\mapsto s_{\omega}(z,z^{\prime})$ is $4L$-Lipschitz for any $z,z^{\prime}\in\mathcal{Z}$. We can directly apply Theorem E.2 with $k=2$ and $p=1$ to $\mathcal{S}$ and get the following maximal inequality:
@@
We obtain an upper-bound on $\mathbb{E}_{\mathbb{D}}{\left[\sup_{\omega\in\Omega}{\left\lvert\tau_{\omega}\right\rvert}\right]}$ by combining the above inequality with Equation 24, then noticing that $2M\leq 2c(\Omega)\max{\left(L\operatorname{diam}(\Omega),M\right)}$ so that:
@@
Appendix F Differentiability Results
The proofs of Propositions A.2 and A.1 are direct applications of the following more general result.
Proposition F.1.

### v3

@@ Appendix F Maximal Inequalities for Bounded and Lipschitz Family of Functions
For a given probability measure $\mu$ on $\mathcal{Z}$, denote by $L_{2}(\mu)$ the space of square $\mu$-integrable real-valued functions. We denote by $\left\lVert f\right\rVert_{\mathbb{D},2}\coloneqq\mathbb{E}_{\mathbb{D}}\left[f(z)^{2}\right]^{\frac{1}{2}}$ the $L_{2}(\mu)$-norm of any function $f\in L_{2}(\mu)$. For any $\epsilon>0$, we denote by $D\left(\epsilon,\mathcal{T},L_{2}(\mu)\right)$ the $\epsilon$-packing number of $\mathcal{T}$ w.r.t. $L_{2}(\mu)$. The next proposition provides a control on such a number under regularity conditions on the family $\mathcal{T}$.
Proposition F.1 (Control on the packing number).
Assume that $\Omega$ is a compact subset of $\mathbb{R}^{d}$, that the parametric family $\mathcal{T}$ defined in Equation˜18 is uniformly bounded by a positive constant $M$, and that there exists a positive constant $L$ so that, for any probability measure $\mu$ on $\mathcal{Z}$, $\omega\mapsto t_{\omega}(z)$ is $L$-Lipschitz for any $z\in\mathcal{Z}$. Then, there exists a positive constant $c(\Omega)$ greater than $1$ that depends only on $\Omega$ and $d$ so that, for any probability measure $\mu$ on $\mathcal{Z}$, the following bound holds for any $0<\epsilon\leq M$:
@@
First using [Kosorok2008, Lemma 9.18] and [Kosorok2008, Paragraph 8.1.2], we know that the $\epsilon$-packing number $D\left(\epsilon,\mathcal{T},L_{2}(\mu)\right)$ is smaller than the $\frac{\epsilon}{2}$-bracketing number $N_{[]}\left(\frac{\epsilon}{2},\mathcal{T},L_{2}(\mu)\right)$. Hence, we only need to control the bracketing number. To this end, we recall that the function $\omega\mapsto t_{\omega}(z)$ is $L$-Lipschitz for any $z\in\mathcal{Z}$, so that [van2000asymptotic, Example 19.7] ensures the existence of a positive constant $c(\Omega)$ that depends only on $\Omega$ for which the following inequality holds for any $0<\epsilon<L\text{diam}(\Omega)$:
@@
Theorem F.2 (Maximal inequality for degenerate, bounded, and Lipschitz $U$-processes).
Let $k$ be either $1$ or $2$. Consider a parametric family $\mathcal{T}\coloneqq\left\{\otimes^{k}\mathcal{Z}\ni(z_{1},\ldots,z_{k})\mapsto t_{\omega}(z_{1},\ldots,z_{k})\in\mathbb{R}\mid\omega\in\Omega\right\}$ of real-valued functions over $\otimes^{k}\mathcal{Z}$ indexed by a parameter $\omega\in\Omega$, where $\Omega$ is a compact subset of $\mathbb{R}^{d}$. For a given probability distribution $\mathbb{D}$ over $\mathcal{Z}$, assume that all elements $t_{\omega}$ are degenerate w.r.t. $\mathbb{D}$, meaning that:
@@
$U_{n}^{k}t_{\omega}\coloneqq\begin{cases}\displaystyle{\frac{1}{n}\sum_{i=1}^{n}t_{\omega}(z_{i})},&\qquad\text{if}\quad k=1\\ \displaystyle{\frac{1}{n(n-1)}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}t_{\omega}(z_{i},z_{j})},&\qquad\text{if}\quad k=2.\end{cases}$
@@
Maximal inequality for degenerate $U$-processes. We will first apply the general result in [sherman1994maximal, Maximal inequality] which controls $\mathbb{E}_{\mathbb{D}}\left[\sup_{\omega\in\Omega}\left\lvert U_{n}^{k}t_{\omega}\right\rvert\right$] in terms of the packing number of $\mathcal{T}$. First note, by assumption, that the functions $t_{\omega}(z_{1},\ldots,z_{k})$ are uniformly bounded by a positive constant $M$. Therefore, the constant function $T(z_{1},\ldots,z_{k})\coloneqq M$ is an envelope for $\mathcal{T}$, i.e., $T$ satisfies $T(z_{1},\ldots,z_{k})\geq\sup_{\omega\in\Omega}\left\lvert t_{\omega}(z_{1},\ldots,z_{k})\right\rvert$ for any $(z_{1},\ldots,z_{k})\in\otimes^{k}\mathcal{Z}$. The envelope $T$ is, a fortiori, square $\mu$-integrable for any probability measure $\mu$ on $\otimes^{k}\mathcal{Z}$. Hence, we can apply [sherman1994maximal, Maximal inequality] with the choice $T$ for the envelope function and set the integer $m$ appearing in the result to $m=d$ to get the following bound:
@@
where $\Gamma$ is a positive universal constant [^2] The constant $\Gamma$ appearing in [sherman1994maximal, Maximal inequality] depends only on $k$, $p$ and $m$, i.e., $\Gamma\coloneqq g(k,p,m)$. Since, we are only interested in $k\leq 2$ and $p\leq 2$ and $m$ is fixed to $d$, we choose $\Gamma$ to be $\max_{1\leq k,p\leq 2}g(k,p,d)^{\frac{1}{p}}$, so that it is the same in all our cases. that depends only on $d$ and that we choose to be greater than $1$, while $\mu_{n}$ are suitably chosen probability measures on $\otimes^{k}\mathcal{Z}$ that possibly depend on the samples $z_{1},\ldots,z_{n}$ and other random variables, and $\delta_{n}\left\lVert T\right\rVert_{\mu_{n},2}\coloneqq\sup_{\omega\in\Omega}\left\lVert t_{\omega}\right\rVert_{\mu_{n},2}$. Here, the expectation symbol in the right-hand side is over all randomness on which $\mu_{n}$ might depend. Note that the original result in [sherman1994maximal, Maximal inequality] is stated using a slightly different definition of the packing number but which is still equivalent to the statement above in our setting [^3] In [sherman1994maximal, Maximal inequality], the author considers a modified version of the $\epsilon$-packing number (call it $\tilde{D}(\epsilon,\mathcal{T},L_{2}(\mu))$) associated to $L_{2}(\mu)$ but endowed with a normalized version of the standard norm on $L_{2}(\mu)$: $\left\lVert f\right\rVert_{\mu}\coloneqq\frac{\left\lVert f\right\rVert_{\mu,2}}{\left\lVert T\right\rVert_{\mu,2}}$. Both numbers are related by the following identity: $\tilde{D}\left(\epsilon,\mathcal{T},L_{2}(\mu)\right)=D(\epsilon{\left\lVert T\right\rVert_{\mu,2}},\mathcal{T},L_{2}(\mu))$, thus making the statement (19) equivalent to the original statement in [sherman1994maximal, Maximal inequality]. .
In our setting, the envelope function is constant and equal to $M$, and by definition $\delta_{n}\leq 1$. Hence, the inequality in Equation˜19 further becomes:
$\mathbb{E}_{\mathbb{D}}\left[\sup_{\omega\in\Omega}\left\lvert U_{n}^{k}t_{\omega}\right\rvert^{p}\right]^{\frac{1}{p}}\leq n^{-\frac{k}{2}}M\Gamma\mathbb{E}\left[\int_{0}^{1}\left(D\left(\epsilon\left\lVert T\right\rVert_{\mu_{n},2},\mathcal{T},L_{2}(\mu_{n})\right)\right)^{\frac{1}{2dp}}\mathop{}\!\mathrm{d}\epsilon\right].$ (20)
@@
Control on the packing number. We have shown that the constant function $T(z_{1},\ldots,z_{k})\coloneqq M$ is an envelope for $\mathcal{T}$ which is, a fortiori, square $\mu$-integrable for any probability measure $\mu$ with $\left\lVert T\right\rVert_{\mu,2}=M<+\infty$. Moreover, the functions $\omega\mapsto t_{\omega}(z_{1},\ldots,z_{k})$ are $L$-Lipschitz for any $(z_{1},\ldots,z_{k})\in\otimes^{k}\mathcal{Z}$. We can therefore apply Proposition˜F.1 which ensures the existence of a positive constant $c(\Omega)$ greater than $1$ and that depends only on $\Omega$ and $d$ so that the following estimate on the $\epsilon$-packing number of the class $\mathcal{T}$ w.r.t. $L_{2}(\mu)$ holds:
@@
Combining Equation˜21 with Equation˜20 yields:
$\mathbb{E}_{\mathbb{D}}\left[\sup_{\omega\in\Omega}\left\lvert U_{n}^{k}t_{\omega}\right\rvert^{p}\right]^{\frac{1}{p}}$ $\leq n^{-\frac{k}{2}}M\Gamma\mathbb{E}\left[\int_{0}^{1}\left(A\epsilon^{-d}\right)^{\frac{1}{2dp}}\mathop{}\!\mathrm{d}\epsilon\right]=n^{-\frac{k}{2}}M\Gamma A^{\frac{1}{2dp}}\underbrace{\int_{0}^{1}\epsilon^{-\frac{1}{2p}}\mathop{}\!\mathrm{d}\epsilon}_{\leq 2}$
@@
The following two propositions are particular instances of Theorem˜F.2 and will be used to obtain the main bounds.
Proposition F.3 (Maximal inequality for empirical processes).
@@
The upper bound is a direct consequence of Theorem˜F.2. Indeed consider the family $\mathcal{S}$ of functions of the form $s_{\omega}(z)=t_{\omega}(z)-\mathbb{E}_{\bar{z}\sim\mathbb{D}}\left[t_{\omega}(\bar{z})\right]$, for any $z\in\mathcal{Z}$. Then clearly, the process $U_{n}^{1}s_{\omega}\coloneqq\frac{1}{n}\sum_{i=1}^{n}s_{\omega}(z_{i})$ is degenerate of order $k=1$, and the family $\mathcal{S}$ is uniformly bounded by $2M$ and is $2L$-Lipschitz. Hence, by Theorem˜F.2, the following maximal inequality holds:
@@
We get the desired upper bound by redefining $c(\Omega)$ to contain the factor $2$. ∎
Proposition F.4 (Maximal inequality for $U$-processes of order 2).
@@
The proof will proceed by first decomposing $\tau_{\omega}$ into a sum of a degenerate $U$-process and a term of order $\mathcal{O}(\frac{1}{n})$. The maximal inequality for degenerate $U$-processes from [sherman1994maximal] will be employed to obtain the desired bound.
@@
$s_{\omega}(z,z^{\prime})=t_{\omega}(z,z^{\prime})-\mathbb{E}_{\bar{z}\sim\mathbb{D}}\left[t_{\omega}(z,\bar{z})\right]-\mathbb{E}_{\bar{z}\sim\mathbb{D}}\left[t_{\omega}(\bar{z},z^{\prime})\right]+\mathbb{E}_{\bar{z},\underline{z}\sim\mathbb{D}\otimes\mathbb{D}}\left[t_{\omega}(\bar{z},\underline{z})\right].$ (22)
@@
$\tau_{\omega}=\overbrace{\frac{1}{n(n-1)}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}s_{\omega}\big(z_{i},z_{j}\big)}^{U_{n}^{2}s_{\omega}}-\frac{1}{n^{2}(n-1)}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}t_{\omega}(z_{i},z_{j})+\frac{1}{n^{2}}\sum_{i=1}^{n}t_{\omega}(z_{i},z_{i}).$
@@
Hence, it only remains to control the first term in the above inequality. To this end, we will use a maximal inequality for degenerate $U$-processes due to [sherman1994maximal].
Maximal inequality for degenerate $U$-processes. We will first check that $U_{n}^{2}s_{\omega}$ is a degenerate statistic for a given $\omega\in\Omega$. Simple calculations show that for any $z$ in $\mathcal{Z}$:
@@
The above equalities precisely ensure that $U_{n}^{2}s_{\omega}$ is a degenerate $U$-statistic for $\mathbb{D}$. Consider now the family $\mathcal{S}\coloneqq\left\{\mathcal{Z}\times\mathcal{Z}\ni(z,z^{\prime})\mapsto s_{\omega}(z,z^{\prime})\in\mathbb{R}\mid\omega\in\Omega\right\}$. We show that $\mathcal{S}$ is uniformly bounded and Lipschitz which allows to directly apply the result stated in Theorem˜F.2, which is a special case of the more general result in [sherman1994maximal, Maximal inequality]. First note, by assumption, that the functions $t_{\omega}(z,z^{\prime})$ are uniformly bounded by a positive constant $M$. Hence, using Equation˜22, it follows that $s_{\omega}(z,z^{\prime})$ is uniformly bounded by $4M$. Moreover, the functions $\omega\mapsto t_{\omega}(z,z^{\prime})$ are $L$-Lipschitz for any $z,z^{\prime}$ in $\mathcal{Z}$. Hence, from Equation˜22, we directly have that $\omega\mapsto s_{\omega}(z,z^{\prime})$ is $4L$-Lipschitz for any $z,z^{\prime}\in\mathcal{Z}$. We can directly apply Theorem˜F.2 with $k=2$ and $p=1$ to $\mathcal{S}$ and get the following maximal inequality:
@@
We obtain an upper bound on $\mathbb{E}_{\mathbb{D}}\left[\sup_{\omega\in\Omega}\left\lvert\tau_{\omega}\right\rvert\right$] by combining the above inequality with Equation˜23, then noticing that $2M\leq 2c(\Omega)\max\left(L\operatorname{diam}(\Omega),M\right)$ so that:
@@
Appendix G Differentiability Results
The proofs of Propositions˜B.2 and B.1 are direct applications of the following more general result.
Proposition G.1.